OpenStudy (anonymous):

MEDALS ALL AROUND... Hey guys, need some steps and some help on how to solve this: (2x-3)/4 + (x-1)/3 =2

OpenStudy (anonymous):

(2x-3)/4 + (x-1)/3 = 2 [3(2x-3) + 4(x-1)]/12 = 2 (6x - 9 + 4x - 4) = 2 * 12 10x - 13 = 24 10x = 37 x = 37/10 = 3.7 Hopefully that helps.

OpenStudy (campbell_st):

multiply each term by 12 common denominator 3(3x -3) + 4(x -1) = 24 expand 3x - 9 + 4x - 3 = 24 collect like terms 7x - 12 = 24 I'll let you solve it from here

OpenStudy (campbell_st):

oops 7x - 13 = 24